# TabMGP: Martingale Posterior with TabPFN

### v1

@@ Appendix H Coverage rate of the marginal credible interval
Table 11: Coverage and Winkler score (median over 100 repetitions, shown in parantheses) of the marginal credible intervals for multinomial logistic regression with the ‘yeast’ data. The method with the lowest Winkler score is shown in bold. The target coverage is 0.95.
$z_{1:n}$ Feature Name TabMGP $B$ Bayes
wine intercept_3 0.97 (2.87) 0.42 (12.1) 0.47 (6.82)
intercept_4 0.99 (2.90) 0.37 (18.9) 0.41 (12.2)
intercept_5 0.98 (3.00) 0.53 (8.18) 0.56 (4.96)
intercept_6 0.99 (3.97) 0.81 (9.82) 0.89 (7.39)
V1_3 0.97 (1.23) 0.72 (2.49) 0.89 (2.35)
V1_4 0.90 (1.31) 0.74 (2.39) 0.85 (2.35)
V1_5 0.88 (1.41) 0.74 (2.51) 0.89 (2.50)
V1_6 0.91 (1.81) 0.74 (3.37) 0.84 (3.55)
V2_3 0.90 (1.16) 0.79 (2.14) 0.89 (2.04)
V2_4 0.81 (1.43) 0.73 (2.25) 0.88 (2.09)
V2_5 0.82 (1.62) 0.69 (2.42) 0.87 (2.36)
V2_6 0.84 (2.12) 0.67 (3.61) 0.80 (3.39)
V3_3 0.96 (1.15) 0.76 (2.62) 0.90 (2.43)
V3_4 0.97 (1.20) 0.72 (2.63) 0.90 (2.43)
V3_5 0.99 (1.37) 0.80 (2.74) 0.90 (2.67)
V3_6 1.00 (1.85) 0.74 (3.62) 0.89 (3.79)
V4_3 0.98 (1.50) 0.70 (2.95) 0.88 (3.06)
V4_4 0.96 (1.55) 0.74 (2.94) 0.91 (2.99)
V4_5 0.93 (1.77) 0.72 (3.07) 0.89 (3.21)
V4_6 0.90 (2.19) 0.71 (4.52) 0.84 (4.32)
V5_3 0.98 (1.34) 0.68 (2.77) 0.90 (3.06)
V5_4 0.96 (1.39) 0.77 (3.00) 0.88 (3.11)
V5_5 0.94 (1.99) 0.82 (3.26) 0.93 (3.68)
V5_6 0.97 (2.49) 0.74 (5.15) 0.90 (5.41)
V6_3 0.79 (2.07) 0.69 (3.86) 0.69 (3.05)
V6_4 0.84 (2.04) 0.70 (3.73) 0.66 (3.08)
V6_5 0.84 (2.19) 0.72 (3.78) 0.68 (3.16)
V6_6 0.83 (2.72) 0.70 (4.75) 0.76 (3.89)
V9_3 0.97 (1.31) 0.75 (2.89) 0.84 (2.65)
V9_4 0.98 (1.35) 0.72 (2.96) 0.84 (2.63)
V9_5 0.95 (1.48) 0.71 (3.05) 0.83 (2.76)
V9_6 0.94 (1.82) 0.73 (3.85) 0.81 (3.74)
V10_3 0.97 (1.25) 0.75 (2.47) 0.86 (2.54)
V10_4 0.96 (1.28) 0.71 (2.63) 0.92 (2.54)
V10_5 0.96 (1.38) 0.74 (2.54) 0.87 (2.63)
V10_6 0.96 (1.64) 0.79 (3.44) 0.87 (3.33)
V11_3 0.93 (1.95) 0.77 (3.01) 0.93 (3.22)
V11_4 0.97 (1.99) 0.73 (3.24) 0.89 (3.24)
V11_5 0.93 (2.17) 0.71 (3.29) 0.88 (3.37)
V11_6 0.93 (2.88) 0.64 (4.51) 0.76 (4.47)
@@
Table 12: Coverage and Winkler score (median over 100 repetitions, shown in parantheses) of the marginal credible intervals for multinomial logistic regression with the ‘wine data. The method with the lowest Winkler score is shown in bold. The target coverage is 0.95.

### v2

@@ Appendix H Coverage rate of the marginal credible interval
Table 11: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for multinomial logistic regression with the ‘yeast’ data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Bayes
Setup Feature Name Rate Size Rate Size Rate Size
alm_ME2 0.93 3.43 0.47 3.86 0.45 2.58
alm_ME3 0.95 3.16 0.56 4.22 0.70 2.81
alm_MIT 0.85 1.20 0.89 1.49 0.71 1.03
alm_NUC 0.93 0.79 0.90 0.97 0.90 0.79
gvh_ME2 0.94 2.39 0.56 3.27 0.57 2.19
gvh_ME3 0.94 1.41 0.82 1.95 0.73 1.49
gvh_MIT 0.92 0.94 0.86 1.20 0.77 0.87
gvh_NUC 0.94 0.61 0.93 0.86 0.86 0.64
intercept_ME2 0.89 6.35 0.39 10.10 0.43 4.42
intercept_ME3 0.95 3.53 0.54 4.87 0.67 3.01
intercept_MIT 0.97 1.08 0.87 1.30 0.82 0.94
intercept_NUC 1.00 0.70 1.00 0.81 1.00 0.62
mcg_ME2 0.90 2.74 0.61 4.14 0.59 2.47
mcg_ME3 0.92 1.63 0.82 2.11 0.81 1.64
mcg_MIT 0.90 0.90 0.90 1.12 0.78 0.85
mcg_NUC 0.96 0.66 0.94 0.89 0.86 0.64
mit_ME2 0.93 2.48 0.57 2.93 0.62 2.62
mit_ME3 0.97 1.99 0.84 2.65 0.77 2.07
mit_MIT 0.87 0.92 0.84 1.07 0.80 0.75
mit_NUC 0.93 0.76 0.90 0.93 0.87 0.70
nuc_ME2 0.98 3.30 0.55 3.92 0.65 3.26
nuc_ME3 0.74 1.67 0.74 2.23 0.69 1.77
nuc_MIT 0.94 1.34 0.83 1.61 0.76 1.28
nuc_NUC 0.89 0.81 0.81 1.04 0.72 0.70
vac_ME2 0.99 2.06 0.63 2.62 0.71 2.31
vac_ME3 0.97 1.60 0.87 2.22 0.91 1.98
vac_MIT 0.98 0.76 0.90 1.04 0.85 0.78
yeast vac_NUC 0.95 0.53 0.89 0.75 0.86 0.57
@@
Table 12: Coverage (rate) and median width (size) over 100 repetitions of marginal credible intervals for multinomial logistic regression with the ‘wine’ data. The method achieving at least the target coverage with the smallest size is shown in bold. If none of the methods achieve the target coverage, the method with the highest coverage is shown in bold. The target coverage is 0.95.
TabMGP $B$ Bayes
Setup Feature Name Rate Size Rate Size Rate Size
V10_3 0.97 1.25 0.75 2.30 0.73 1.65
V10_4 0.96 1.27 0.71 2.18 0.71 1.61
V10_5 0.96 1.34 0.74 2.25 0.70 1.69
V10_6 0.96 1.64 0.79 3.03 0.73 2.09
V11_3 0.93 1.94 0.77 2.72 0.76 2.09
V11_4 0.97 1.98 0.73 2.78 0.78 2.09
V11_5 0.93 2.16 0.71 2.93 0.78 2.21
V11_6 0.93 2.83 0.64 4.22 0.67 2.85
V1_3 0.97 1.21 0.72 2.22 0.74 1.50
V1_4 0.90 1.26 0.74 2.20 0.72 1.48
V1_5 0.88 1.35 0.74 2.33 0.68 1.59
V1_6 0.91 1.73 0.74 3.17 0.74 2.19
V2_3 0.90 1.12 0.79 1.97 0.74 1.29
V2_4 0.81 1.30 0.73 2.06 0.68 1.36
V2_5 0.82 1.45 0.69 2.23 0.70 1.51
V2_6 0.84 1.99 0.67 3.12 0.66 2.15
V3_3 0.96 1.14 0.76 2.16 0.65 1.49
V3_4 0.97 1.18 0.72 2.21 0.63 1.49
V3_5 0.99 1.37 0.80 2.33 0.76 1.66
V3_6 1.00 1.85 0.74 3.19 0.71 2.37
V4_3 0.98 1.49 0.70 2.60 0.74 1.94
V4_4 0.96 1.50 0.74 2.66 0.75 1.93
V4_5 0.93 1.65 0.72 2.69 0.75 2.04
V4_6 0.90 2.17 0.71 3.78 0.64 2.66
V5_3 0.98 1.33 0.68 2.46 0.74 1.74
V5_4 0.96 1.37 0.77 2.55 0.77 1.77
V5_5 0.94 1.98 0.82 2.90 0.81 2.22
V5_6 0.97 2.47 0.74 4.78 0.75 3.36
V6_3 0.79 1.77 0.69 3.10 0.56 1.75
V6_4 0.84 1.78 0.70 3.06 0.57 1.77
V6_5 0.84 1.92 0.72 3.19 0.60 1.85
V6_6 0.83 2.15 0.70 4.07 0.59 2.37
V9_3 0.97 1.29 0.75 2.50 0.64 1.72
V9_4 0.98 1.34 0.72 2.50 0.65 1.74
V9_5 0.95 1.46 0.71 2.57 0.65 1.82
V9_6 0.94 1.81 0.73 3.51 0.70 2.33
intercept_3 0.97 2.87 0.42 5.76 0.55 2.76
intercept_4 0.99 2.90 0.37 5.72 0.49 2.75
intercept_5 0.98 3.00 0.53 5.91 0.56 2.82
wine intercept_6 0.99 3.94 0.81 9.06 0.69 4.35